OpenStudy (anonymous):

A parallelogram with sides of 6 and 10 has an area of 30. Find the measure of each angle of the parallelogram.

OpenStudy (btaylor):

 In the diagram, sin(x) = 3/6, so x = arcsin(1/2) = 30º The other angle is supplementary, so it is 150º